OpenStudy (anonymous):

Find the value of c that makes each trinomial a perfect square x^2-18x+c

OpenStudy (anonymous):

(x+9)^2=x^2-18x+81 therefore c=81

OpenStudy (anonymous):

where (x+9)^2 is a perfect square